hartnn (hartnn):

Prove that f_1 (x)=1,f_2 (x)=x,f_3 (x)=((3x^2-1))/2 are orthogonal over (-1,1) I know the method, Just need a start general form for \(f_m (x) ~~ and~~ f_n(x)\)

hartnn (hartnn):

\( f_1 (x)=1, \\ f_2 (x)=x,\\ f_3 (x)=((3x^2-1))/2 \)

OpenStudy (math&ing001):

Isn't this how you do it ? \[\int\limits_{-1}^{1}x.\frac{ 3x ^{2}-1 }{ 2 }=\int\limits_{-1}^{1}(3/2 .x ^{3}-x/2)dx = 0\]

hartnn (hartnn):

for m\(\ne\) n, \( ∫_a^b [f_n (x)] [ f_m (x)] dx=0\) for \(m= n\) \(∫_a^b [f_n (x)]^2 dx \ne 0\) this is how i do it

OpenStudy (math&ing001):

So what's the general form ?

ganeshie8 (ganeshie8):

just need to work 3 integrals, right ? m,n : 1,2 1,3 2,3

ganeshie8 (ganeshie8):

how do you put them in matrix @OOOPS ?

hartnn (hartnn):

need to prove 1,1 2,2 3,3 not =0 also, right ? or not required ?

OpenStudy (anonymous):

A true scale plot is attached.

hartnn (hartnn):

1,3 is not an odd function, i doubt it would be 0 ... 1,2 and 2,3 are odd, directly 0

hartnn (hartnn):

oh yeah, all 3 are 0

ganeshie8 (ganeshie8):

discrete : \(\large a.b = 0\) means \(a\) and \(b\) are orthogonal so i think proving first condition is sufficient, not sure though.

hartnn (hartnn):

ok, i am quite sure i need to prove 1,1 ...2,2...3,3 not =0 too will do it

ganeshie8 (ganeshie8):

continuous : \(\large \int a.b = 0 \implies a, b \) are orthogonal and \(0\) vector is orthogonal to every vector, so im not sure why you're quite sure haha :3

hartnn (hartnn):

wait, why ?

hartnn (hartnn):

lol, i didn't get ganeshie's point :P

hartnn (hartnn):

and 1,1 will be 0 :O

OpenStudy (math&ing001):

Actually 1,1 can't be =0

hartnn (hartnn):

oh yes, its not odd function ok so last doubt, why its not requied to show 1,1 2,2 3,3 not =0

OpenStudy (math&ing001):

And you're right 2.2 and 3.3 shouldn't =0 either.